OpenStudy (anonymous):

Solve the quadratic equation by completing the square. x^2+16x+57=0

OpenStudy (anonymous):

First, choose the appropriate form and fill in the blanks with the correct numbers. Then, solve the equation. Round your answer to the nearest hundredth. If there is more than one solution, separate them with commas. FORMS: A. (x+?)^2 = 0 B. (x-?)^2 = 0

OpenStudy (triciaal):

do you know what is meant by a "perfect square?"

OpenStudy (anonymous):

It doesn't say perfect square.

OpenStudy (triciaal):

I know that is the idea (a + b)^2 = (a + b)(a + b) = a^2 + 2ab + b^2

OpenStudy (triciaal):

you want to get in this format and that is why it is called "completing-the-square"

OpenStudy (triciaal):

here are some basic steps with the equation in standard form a x^2 + bx + c = 0 divide everything by a subtract c find 1/2 of b then square (x +b) notice you will have b^2 added on the left so subtract and keep the balance (or you can add to the other side in one step) you should have (x + b/2)^2 = -c + (b/2)^2

OpenStudy (triciaal):

do you understand so far? can you continue?

OpenStudy (triciaal):